OpenStudy (anonymous):

**HELP NEEDED PLEASE!!* The Richter Scale is used to measure the magnitude of an earthquake. The magnitude, (R), is given by the equation R=0.65log(0.39E) +1.45, where E is the energy released by the earthquake. Determine the magnitude of 7.5.

OpenStudy (anonymous):

Simplify so that the answer is written in form b=log a

OpenStudy (ganpat):

U mean, Determine energy released ??

OpenStudy (anonymous):

I'm not sure, the question said determine the magnitude, but it's probably a mistake

OpenStudy (anonymous):

I got the answer of E = 5207481079.2429121, but I just don't no how to write it in form b=loga

OpenStudy (ganpat):

ok, if u can use the calculator, Than calculate for E by substituting R in above given equation..

OpenStudy (anonymous):

I found E, but I just don't know how to write it in proper form, b=loga

OpenStudy (ganpat):

7.5/0.65 = log (0.39E) 11.53 = log (0.39E)... does it make sense :-p

OpenStudy (anonymous):

yes thank youu. so would that be the final answer?

OpenStudy (anonymous):

wait, what about the 1.45?

OpenStudy (ganpat):

i guess so, substituting value of E will extend beyond limit.. or you can try to bring out 0.39 out of it ..

OpenStudy (ganpat):

bring it on left hand side and subtract to 11.53

OpenStudy (ganpat):

i missed that :D, i miss a lot man..

OpenStudy (anonymous):

lol. thanks so much! you practically just saved my life. haha like not even kidding.

OpenStudy (ganpat):

11.53 = log (0.39E) + 1.45 11.53 - 1.45 = log (0.39E) 10.08 = log (0.39E)

OpenStudy (ganpat):

my pleasure bro !!! :)

OpenStudy (anonymous):

***girl :)

OpenStudy (ganpat):

oopsss. sorry !!

OpenStudy (anonymous):

its all good